OpenStudy (anonymous):

Look at the numbers below. 5, 6, 9, 11 Which pair of numbers has an even product and an odd sum? A. 5 and 9 B. 5 and 11 C. 6 and 9 D. 9 and 11

OpenStudy (ash2326):

Kyrie try yourself this one, if you get stuck, we're here to help you.

OpenStudy (anonymous):

But Idk where to even begin! D:

OpenStudy (ash2326):

Remember that two odd no.s always add up to even

OpenStudy (anonymous):

well yes but.....

OpenStudy (anonymous):

But 6 is the only even....the rest are odd....

OpenStudy (ash2326):

yeah, where you are stuck, check out the options one by one. I'll do one then you try the other options 5 and 9 5*9=45 it's odd, we want product to be even 5+9= 14 it's even, we want sum to be odd so this is not our answer, try other options.

OpenStudy (ragingsquirrel):

C...

OpenStudy (anonymous):

I didnt know wether to add,subtract,multiply,or devide...idk :((((

OpenStudy (ragingsquirrel):

First, you multiply the numbers to see if they come out even. If some do, you add the numbers together to see if they come out odd.

OpenStudy (ash2326):

sum is add product is multiplication now you try

OpenStudy (ash2326):

Kyrie if you try you'll learn, don't you want to be a great mathematician ?

OpenStudy (anonymous):

5*11=55...odd 5+11=16..even....So no this isn't the answer

OpenStudy (ash2326):

yeah great going:)

OpenStudy (anonymous):

so like 6*9=54...even 6+9=15...odd So C is the answer Right? So Alex was right..

OpenStudy (ash2326):

yeah great Kyrie:)

OpenStudy (anonymous):

bc 9*11=99...odd 9+11=20...even So it has to be C. So THank you Alex and Ash very much!(:

OpenStudy (ash2326):

welcome Kyrie:) Always try, you'll learn more and if you get stuck we're here to help you.